integrate [x^2]dx , limit of integration= 0 to 1.5

Grade:12th pass

4 Answers

Sunil Kumar FP
10 years ago
Between0and1the function is0, integral0. Between1and2v, the function is1, integral(1)(2v-1). Between2vand1.5, the function is2, integral(2)(1.5-2v). Add up and simplify. I think you will get2-2v

thanks and regards
sunil kr
askIItian faculty
chaitnyakishore
10 years ago
integration 0 to 1.5 (x^2)=[x^3/3]limits 0 to 1.5 =(1.5^3/3)-(0^3/3) =(3.375/3)-(0) =1.125
chaitnyakishore
10 years ago
integration 0 to 1.5 (x^2)dx=[x^3/3]limits 0 to 1.5 =(1.5^3/3)-(0^3/3) =(3.375/3)-(0) =1.125
VEERU YERRAYYA
10 years ago
1.5 1.5 ? x².dx = x³/3 ] 0 (1/3)(1.5³-0³)=1.125